# arXiv:2603.00183v1[cs.SE] 26 Feb 2026

### v1

@@ 6.1 Threats to validity
Internal validity: The main threat to internal validity concerns the correctness of the implementation of the proposed solution. Multiple tactics were used to reduce this risk. The entire contributed project uses typechecking and linting, with all rules enabled by default. The proposed framework has correctness precautions built in. Finally, the research artifacts are published for external inspection. In terms of SR, the number of involved researchers poses the primary threat to its internal validity. To address this threat, state-of-the-art guidelines Wohlin [2014], Moher et al. [2015], Kitchenham et al. [2023] were used, a SR protocol was written, and then approved by the second author. The use of the snowballing process should also reduce internal validity threats.
External validity: The main risk of external validity comes from the fact that only open-source Java projects were tested. The reached conclusions might not generalize to projects written in other languages or to proprietary solutions. We believe that despite the reported threats, the use of RTPTorrent was the correct choice for the study. The dataset contains real-world data; it includes projects with diverse maturities, contributor counts, test case counts, codebase sizes, and failure percentages. For the SR, external validity is slightly concerned with the choice to select only peer-reviewed studies written in English. However, the vast majority of articles related to TCP are written in English. Furthermore, due to the use of Google Scholar, the external validity threats resulting from publisher and database bias are greatly reduced Wohlin [2014].
Construct validity: The primary risks in this area refer to the use of less popular $r\text{APFD}$ instead of APFD, and the use of newly proposed $r\text{APFD}_{C}$ and ATR metrics. The threat arising from the usage of $r\text{APFD}$ and $r\text{APFD}_{C}$ is considered minimal. Most importantly, these metrics have the same monotonicity as APFD and $\text{APFD}_{C}$, respectively. Thus, it does not harm the relative comparisons of different algorithms. In terms of absolute evaluation, the rectified metrics have much more convenient semantics. For the proposed ATR metric, although it has a logical interpretation, it is very different from NTR. To eliminate associated risks, all applicability results are reported in terms of both metrics. Furthermore, we avoid invalid measurement practices reported in the literature Bagherzadeh et al. [2022], Zhao et al. [2023].
@@ 6.2 Recommendations for researchers and practitioners
Validity of TCP research: We share our insights to increase the validity of the TCP research ecosystem. Firstly, following Mattis et al. [2020], we observe that TCP studies are often performed on proprietary data or projects that do not reflect real-world software development practices. Secondly, not all studies are reproducible, and some do not publish their results at all Catal and Mishra [2013]. Finally, incorrect use of metrics occurs in the literature Bagherzadeh et al. [2022], Zhao et al. [2023]. We suggest adhering to the following recommendations. Studies should use open-source datasets with diverse subject programs to support reproducible research and increase ecological validity. Their results should be compared with baselines better than random prioritization. Lastly, researchers should be careful when calculating evaluation metrics; in particular, we propose using rectified metrics from the APFD family and averaging over only failed cycles.
@@
Approach combinators: There are areas for improvement of the methodology. Firstly, additional combinators and categories may be proposed. Although combinators are applicable to almost all TCP methods, we only combined heuristic methods. Furthermore, we see an option to use supervised ML strategies on some combinators; namely, the weights of mixers may be automatically learned.
@@ 7 Conclusions
The literature survey conducted led to a few key insights. There are many TCP evaluation tools available, and none of them are universally used. The choice of the dataset should be influenced by the exact problem statement and other factors. For general use, we recommend the RTPTorrent Mattis et al. [2020] dataset, since it is publicly available and includes diverse subject programs. It is hard to establish the state-of-the-art TCP algorithms as studies from this area are mostly incomparable due to divergences in datasets, metrics, tools, information factors, and problem statements used. Some studies do not empirically evaluate their proposed approaches at all. The most widely used evaluation reference points are base order, random order, and the ROCKET Marijan et al. [2013] algorithm.
@@
Reproducible research (RR) can address problems with the validity of findings in software engineering research Madeyski and Kitchenham [2017]. To support RR and good scientific practices, the research is backed by project files in the form of code, scripts, and SR artifacts. The complete source code and instructions for obtaining and preparing the dataset are available at https://github.com/LechMadeyski/MSc25TomaszChojnacki. See also Section 2 of our online Appendix for details on running the code and extending the project3.

### v2

@@ 6.1 Threats to validity
Internal validity: The main threat to internal validity concerns the correctness of the implementation of the proposed solution. Multiple tactics were used to reduce this risk. The entire contributed project uses typechecking and linting, with all rules enabled by default. The proposed framework has correctness precautions built in. Finally, the research artifacts are published for external inspection. In terms of SR, the number of involved researchers poses the primary threat to its internal validity. To address this threat, state-of-the-art guidelines (Wohlin, 2014; Moher et al., 2015; Kitchenham et al., 2023) were used, a SR protocol was written, and then approved by the second author. The use of the snowballing process should also reduce internal validity threats.
External validity: The main risk of external validity comes from the fact that only open-source Java projects were tested. The reached conclusions might not generalize to projects written in other languages or to proprietary solutions. We believe that despite the reported threats, the use of RTPTorrent was the correct choice for the study. The dataset contains real-world data; it includes projects with diverse maturities, contributor counts, test case counts, codebase sizes, and failure percentages. For the SR, external validity is slightly concerned with the choice to select only peer-reviewed studies written in English. However, the vast majority of articles related to TCP are written in English. Furthermore, due to the use of Google Scholar, the external validity threats resulting from publisher and database bias are greatly reduced (Wohlin, 2014).
Construct validity: The primary risks in this area refer to the use of less popular $r\text{APFD}$ instead of APFD, and the use of newly proposed $r\text{APFD}_{C}$ and ATR metrics. The threat arising from the usage of $r\text{APFD}$ and $r\text{APFD}_{C}$ is considered minimal. Most importantly, these metrics have the same monotonicity as APFD and $\text{APFD}_{C}$, respectively. Thus, it does not harm the relative comparisons of different algorithms. In terms of absolute evaluation, the rectified metrics have much more convenient semantics. For the proposed ATR metric, although it has a logical interpretation, it is very different from NTR. To eliminate associated risks, all applicability results are reported in terms of both metrics. Furthermore, we avoid invalid measurement practices reported in the literature (Bagherzadeh et al., 2022; Zhao et al., 2023).
@@ 6.2 Recommendations for researchers and practitioners
Validity of TCP research: We share our insights to increase the validity of the TCP research ecosystem. Firstly, following Mattis et al. (2020), we observe that TCP studies are often performed on proprietary data or projects that do not reflect real-world software development practices. Secondly, not all studies are reproducible, and some do not publish their results at all (Catal and Mishra, 2013). Finally, incorrect use of metrics occurs in the literature (Bagherzadeh et al., 2022; Zhao et al., 2023). We suggest adhering to the following recommendations. Studies should use open-source datasets with diverse subject programs to support reproducible research and increase ecological validity. Their results should be compared with baselines better than random prioritization. Lastly, researchers should be careful when calculating evaluation metrics; in particular, we propose using rectified metrics from the APFD family and averaging over only failed cycles.
@@
Approach combinators: There are areas for improvement of the methodology. Firstly, additional combinators and categories may be proposed. Although combinators are applicable to almost all TCP methods, we only combined heuristic methods. Furthermore, we see an option to use supervised ML strategies on some combinators; namely, the weights of mixers may be automatically learned. Based on our empirical observations, we offer guidelines for effective combinator construction: (1) combine methods exploiting complementary information sources, as the best-performing combinators in our study integrated orthogonal factors; (2) use balanced weights unless empirically justified otherwise; (3) prefer theoretically grounded aggregation methods such as Borda count over random selection. Configurations combining redundant methods or using extreme weight distributions should be avoided.
@@ 7 Conclusions
The literature survey conducted led to a few key insights. There are many TCP evaluation tools available, and none of them are universally used. The choice of the dataset should be influenced by the exact problem statement and other factors. For general use, we recommend the RTPTorrent (Mattis et al., 2020) dataset, since it is publicly available and includes diverse subject programs. It is hard to establish the state-of-the-art TCP algorithms as studies from this area are mostly incomparable due to divergences in datasets, metrics, tools, information factors, and problem statements used. Some studies do not empirically evaluate their proposed approaches at all. The most widely used evaluation reference points are base order, random order, and the ROCKET (Marijan et al., 2013) algorithm.
@@
Reproducible research (RR) can address problems with the validity of findings in software engineering research (Madeyski and Kitchenham, 2017). To support RR and good scientific practices, the research is backed by project files in the form of code, scripts, and SR artifacts. The complete source code and instructions for obtaining and preparing the dataset are available online9. See also Section 2 of our online Appendix for details on running the code and extending the project1.